JAHN-TELLER AND SPIN-ORBIT COUPLING IN THE ^{2}E GROUND STATE OF CF_{3}O AND CF_{3}S

\underline{TIMOTHY A. BARCKHOLTZ} AND TERRY A. MILLER , Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

The electronic spectra of the Ã ^{2}A ↔ X̃ ^{2}E transition of the CF_{3}O and CF_{3}S radicals are complex because of Jahn-Teller activity in the ^{2}E ground states. Proper analyses of the spectra require the inclusion of both Jahn-Teller coupling and spin-orbit coupling. For CF_{3}O, the couplings are similar in magnitude whereas in CF_{3}S the spin-orbit coupling clearly dominates, although Jahn-Teller coupling cannot be ignored. The analysis of the emission spectra have included first-order and second-order Jahn-Teller coupling constants in each of the three Jahn-Teller active e modes, as well as spin-orbit coupling. Results of our analyses will be presented along with improvements to the computational program.